MOLECULAR ROTATION IN QUANTUM LIQUIDS

\underline{B. SARTAKOV}, General Physics Institute RAS, Vavilov str. 38, 119991 Moscow, Russia; A. VILESOV, Department of Chemistry, University of Southern California, Los Angeles, CA 90089, USA; and J. PETER TOENNIES, Max-Planck-Institut fur Stromungsforschung, Bunsenstrasse 10 37073 Goettingen, Germany.

Rotationally resolved IR spectra of molecules and hydrogen clusters trapped in ^{4}He and mixed ^{3}He/^{4}He droplets are presented. The newest analysis of spectra in terms of effective rotational constants of OCS-(pH_{2})_{n} and OCS-(oD_{2})_{n} complexes with n=11,…16 is presented. Whereas molecular rotation in ^{4}He can be modelled by a free rotor with effective rotational constants the effect of interaction between a molecule and its H_{2} solvation shell is more pronounced. Analysis of the spectra within the frameworks of a rigid symmetric H_{2} shell proves the inconsistency of the rigid shell model with the experimental observations. Furthermore it suggests existence of some low energy excitations (E∼0.1 cm^{−1}) which are tentatively assigned to tunnelling states of the H_{2} cluster.